OpenStudy (anonymous):

If y varies directly as x, and y=5 when x=8, find y when x=4

OpenStudy (anonymous):

sooo, \[5=k(8) \] and \[y=k(4)\]

OpenStudy (anonymous):

\[k = 5/8\] so \[y = (5/8)(4)\]

OpenStudy (anonymous):

Ooh, I see what you did ! Thank you, this is much appreciated.